How would I start to solve this problem? A plane leaves an airport at 3 pm. At 4 pm. another plane leaves the same airport traveling the same direction at a speed 150 miles per hour faster than the first plane. Four hours after the first plane takes off, the second plane is 250 miles ahead of the first plane. Find the speed of each plane. How far did the second plane travel?
Answer by khwang(438)

Sol: Let the speed of the first plane be x miles/hr and then the speed of the 2nd one is x+150 mile/hr.
At 7pm(4 hours after the 1st plane took off) the 2nd plane flew 3hrs and was
250 miles ahead of the 1st plane.
So we have 3(x+150) = 4x +250,
Or 3x + 450 = 4x + 250, So, x = 200 miles/hr and x + 150 = 350 ,
Hence,the speed of the 2nd plane is 350 miles/hr and it travelled
3*350 = 1050 miles.